# Coherent Source Subsampling: A Data-Driven Strategy for Restoring Causal–Acausal Symmetry in Ambient Seismic Wavefield Correlations

### v1

@@ Abstract
Ambient noise tomography relies on the assumption that the seismic wavefield is equipartitioned, meaning that energy is uniformly distributed among all directions. However, in practice, ambient noise sources are highly non-uniform in both spatial and temporal dimensions, resulting in biased estimation of the Green’s function between stations. We introduce a data-driven method, Coherent Source Subsampling (CSS), which selects and averages only those cross-correlation time windows that are associated with the excitation of sources in the stationary-zone. By confining the ensemble average to these coherent subsets, CSS effectively mitigates the influence of anisotropic or intermittent sources and restores causal–acausal symmetry in the retrieved Green’s functions. Applications to regional-scale ambient noise datasets demonstrate that CSS boosts inter-station coherence and enhances the reliability of surface-wave dispersion measurements, providing a physically interpretable bridge between source statistics and noise correlation theory.
@@ 1 Introduction
Surface-wave tomography has long been a cornerstone of the imaging of the Earth’s crust and upper mantle. The resolution of traditional surface-wave tomography, which is based on seismic waves generated by earthquakes recorded by regional and global seismic networks, is fundamentally constrained by the spatial and temporal distribution of seismicity (trampert2003global; ritzwoller1998eurasian; ritzwoller2001crustal; shapiro2002monte; ekstrom2011global; mordret2013near). Over the last 20 years, ambient seismic noise cross-correlations have supplemented surface wave tomography (shapiro2004emergence; sabra2005surface; yang2007ambient; lin2008surface; yang2008structure; yao2009analysis; ritzwoller2011ambient).
@@
of $M_{ij}$ time windows provides an estimator of expectation $\mathbb{E}[\mathbf{w}_{ij}]$. Many authors have demonstrated that, in theory, the empirical average in Eq. 2 converges to the interstation Green’s function, provided the ambient wavefield is isotropic, uncorrelated, and equipartitioned (campillo2003long; wapenaar2004retrieving; wapenaar2006green; snieder2010imaging). Another important theoretical result is that, when the above conditions are satisfied, $\widehat{\mathbb{E}}\left[{\mathbf{w}}_{ij}\right]$ should exhibit causal–acausal symmetry (wapenaar2004retrieving; snieder2004extracting), representing equivalence between waves traveling in both directions between the receivers. In practice, however, this symmetry is rarely achieved (bensen2007processing; stehly2006study; yao2009analysis; tsai2009establishing) because real ambient noise sources are not isotropic (as shown in Fig. 1a) with seasonal variations in their distribution and bursts of body wave energy (stehly2007traveltime; zhan2013spurious; delaney2017passive).
@@
where $\mathrm{concat}(\cdot)$ denotes the concatenation of two or more vectors, and $\mathrm{rev}(\cdot)$ denotes the reversal of the order of vector elements. Here, $\prescript{+}{}{\mathbf{w}}_{ij}$ and $\prescript{-}{}{\mathbf{w}}_{ij}$ denote the vectors associated with the causal and acausal branches of the cross-correlation, respectively, each arranged in order of increasing lag magnitude. The lack of symmetry between the causal and acausal branches means that $\widehat{\mathbb{E}}\left[\prescript{+}{}{\mathbf{w}}_{ij}\right]\neq\widehat{\mathbb{E}}\left[\prescript{-}{}{\mathbf{w}}_{ij}\right]$. If this is the case, then interstation tomography using either of the branches is not justified, since the assumption of equipartitioning is violated. In such cases, the empirical averages are also sensitive to the source distribution and more complicated sensitivity kernels have to the used during tomography (fichtner2015source; ermert2017ambient), which is computationally expensive.
@@
In this paper, we present a data-driven approach, termed Coherent Source Subsampling (CSS), which subsamples the set $\left\{\mathbf{w}_{ij}^{(m)}\right\}$ based on the state of ambient noise sources. We denote the source state using a discrete variable $\lambda$, and two such source states are illustrated in Fig. 1a, using dotted and dashed lines. Our workflow for a given branch $b\in\{+,-\}$, receiver pair $(i,j)$ and source state $\lambda$ is very simple:
@@
Here, $\left|\cdot\right|$ denotes the cardinality of a set, and note that the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is not only pair specific but also branch specific, meaning the sources associated with a particular source state differ between causal and acausal branches (see Fig. 1a). A straightforward approach to conditional averaging is to partition the data according to observable external variables, such as season, month, or time of day. For example, separate averages for winter and summer months could be computed, expecting coherent cross-correlations due to ocean activity or storms within a season (yang2008characteristics; zhan2013spurious). However, such conditioning strategies may not necessarily align with the actual coherence states in the ambient wavefield.
We adopt the representation learning framework to identify coherent source states in a data-driven manner (as shown in Fig. 1b). Specifically, we use a variational autoencoder to identify the subset of windows associated with a particular source state $\lambda$, rather than on predefined external variables. This idea is demonstrated using a synthetic experiment in the Supplementary Text S1. As an example of the source state, consider the stationary-phase zone for a given branch, where the sources coherently contribute to the inter-station response. In this paper, we label the source state for this zone as $\lambda=1$. Another source state is illustrated in Fig. 1b. We show that CSS effectively restores causal–acausal symmetry when averaging cross-correlations confined to a particular source state. Specifically, we achieve
@@
for a choice of source state $\lambda$ and station pair $(i,j)$. Although the causal and acausal branches are redundant in the sense that they contain the same information about the interstation Green’s function, the achievement of causal–acausal symmetry is important for two reasons. First, it validates the Green’s function approximation, which is a prerequisite for robust dispersion and travel-time measurements for tomography. Second, symmetry ensures that there is no contamination from directional source effects and other transients. To our knowledge, this is the first work that achieves the theoretically predicted high degree of causal–acausal symmetry in ambient noise cross-correlations even when the underlying source distribution is highly non-uniform and time-varying. Although CSS performs nonlinear averaging (similar to e.g., phase-weighted stacking by  schimmel1997noise), it is unique in the sense that it achieves causal–acausal symmetry.
@@ 2.1 Noise crosscorrelations
We observed that the choice of the study region is less important to the CSS framework, which is applicable to any ambient noise dataset. However, to validate the method, we used continuous ambient seismic data from the eastern United States, a stable cratonic region characterized by low to moderate seismicity and relatively homogeneous crustal structure (liang2008ambient). In this case, low-frequency vertical-component (LHZ) displacement data from $10$ stations of the US network (Figure S11) were recorded at a sampling rate of 11\,Hz over the interval from January $2005$ through December $2007$ to analyze Rayleigh-wave propagation. The waveform data were corrected for the instrument response, detrended, demeaned, and decimated to a sampling rate of 0.50.5\,Hz. The noise is segmented into non-overlapping $30$ minute windows with a total $M_{ij}$ time windows for each pair $(i,j)$.
Before cross-correlation, each window was temporally normalized using a running absolute mean. The preprocessed traces were then cross-correlated for all available station pairs using FastXC (wang2025high), with a maximum time lag of +1000+1000 s. The cross-correlations are spectrally whitened in the 0.0090.0090.2490.249\,Hz frequency band to minimize the effects of the source signature. The resulting cross-correlations $\mathbf{w}_{ij}$ were then decomposed into their causal and acausal branches as in Equation 3. Before detailing the network architecture and subsampling procedure, we will first give a simple explanation of stationary zones, highlighting the issue to the empirical average in eq. 2, and motivating the need for the conditional average in eq. 4.
@@ 2.2 Stationary-zone source state
Although the cross-correlations are coherent within stationary zones, in practice, the energy contribution from such zones does not necessarily dominate the ensemble average over all available windows (Eq. 2) — this is because the source density $\rho(\mathbf{s},\mathbf{x}_{s})$ may be negligibly small in those regions. In other words, as Eq. (7) indicates that the resulting distribution $\rho(\tau_{ij},\mathbf{s})$ depends jointly on the mapping $T_{ij}(\mathbf{x}_{s})$, and on the spatial source distribution $\rho(\mathbf{x}_{s},\mathbf{s})$, the departure of $\rho(\mathbf{x}_{s},\mathbf{s})$ from theoretical assumptions biases the interstation Green’s function retrieval. In such cases, the retrieval of Green’s functions can be improved by isolating and averaging only windows from stationary zones. We fix $\lambda=1$ to denote the source state of the stationary zone, which is the only state of interest in this study. Additional source states may exist, namely alternative configurations of the ambient noise field, defined by sources with similar signatures that yield coherent cross-correlations between receiver pairs for a given arrival; we label these by $\lambda=2,\ldots,K$. Fig. 1a and the synthetic experiment (Supplementary Text S11) provide representative examples of such states.
@@ 2.3 Latent variable modeling
To enable data-driven identification of coherent source subsets $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$, we adopt a latent-variable modeling approach based on variational autoencoders (VAEs). VAEs provide a probabilistic framework for learning interpretable latent representations (kingma2013auto; doersch2016tutorial; bishop2023deep; prince2023understanding) from high-dimensional data. Within this framework, we introduce a discrete latent variable $\lambda$ (morton2021scalable) to represent source states and design the network architecture such that $\lambda$ correlates with physically meaningful source states, such as the stationary-zone state. The design is based on deliberate architectural decisions: a permutation-invariant encoder that accumulates coherent information (bharadwaj2024extracting) across all windows for a given pair and a categorical posterior distribution over $\lambda$ that naturally groups coherent cross-correlation windows. We emphasize that alternative network designs may exist; however, the focus of this paper is not to exhaustively explore all possible architectures, but rather to demonstrate that a carefully designed VAE can successfully solve the CSS problem by learning to distinguish stationary-zone windows from incoherent noise. The following subsections brief the generative and inference models that implement this approach.
@@ 2.3.1 Generative model
We begin with a generative model that mathematically describes how observed cross-correlations are synthesized from latent variables. That is to say, we train a decoder network that generates each cross-correlation waveform $\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}$ using three distinct latent-variable components — they are highlighted in Fig. 1b.
@@
Here, $l$ denotes the source state with the highest posterior probability for the window $m$, effectively selecting a single coherent feature $\prescript{b}{l}{\mathbf{g}}_{ij}$ for generation. We implement the $\mathtt{decoder}$ as a convolutional neural network.
@@ 2.3.2 Inference model
Here, $\mathtt{enc}_{\mathbf{g}}$, $\mathtt{enc}_{\pi}$, and $\mathtt{enc}_{\mathbf{n}}$ are convolutional encoder components that infer coherent source features, source state probabilities, and nuisance features, respectively. The symmetric encoder component $\mathtt{enc}_{\mathbf{g}}$ exploits the permutation invariance by accumulating information across all time windows for a given station pair and branch. This component infers $\prescript{b}{\lambda}{\mathbf{g}}_{ij}$, which is constrained to be invariant to window ordering. On the contrary, $\mathtt{enc}_{\pi}$ and $\mathtt{enc}_{\mathbf{n}}$ operate on individual windows to capture window-specific features, inferring $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ and $\prescript{b}{}{\mathbf{n}}^{(m)}_{ij}$, respectively. Note that the autoencoder may assign the stationary-zone state to any index in $\{1,\cdots,K\}$. After training, we therefore relabeled the source states so that the stationary-zone state is always set to $\lambda=1$. Additional architectural and training details are given in the Supplementary Text S22.
@@ 2.4 Coherent source subsampling (CSS)
where the condition isolates contributions from sources associated with the state indexed by $\lambda$. It uses the learned source state probabilities $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ for subsampling. Specifically, for each branch $b$, station pair $(i,j)$, and source state $\lambda$, the CSS estimator of the conditional expectation in Eq. 11 is given by Equation 4, where the subset of correlation windows $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is defined as:
@@
It is important to note that the identification of windows belonging to a specific source state is accomplished by probability thresholding (we used $\alpha=0.85$). By averaging only the time windows with high posterior probability $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$, we in effect subsample the windows associated with a given source state. With this targeted subsampling and by setting $\lambda=1$, we can isolate the contributions from stationary regions in the source field. Note that the conditional average $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]$ in Eq. 4 is a non‑linear function of the data, unlike linear averaging, because the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ itself depends on the data. It is important to emphasize that the averaging is performed on the original cross-correlation windows, using the VAE only to identify which windows belong to each source state. Apart from the source state probabilities, we do not use any other outputs of the encoder or decoder networks, such as the reconstructed cross-correlations, in our analysis. We now demonstrate the advantages of subsampling over indiscriminate averaging.

### v2

@@ Abstract
Ambient noise tomography relies on the assumption that the seismic wavefield is equipartitioned. In practice, ambient noise sources are spatially and temporally heterogeneous, producing biased estimates of Green’s function between stations. We introduce a data-driven method, Coherent Source Subsampling (CSS), which selects and averages only cross-correlation time windows associated with excitation of sources in the stationary zone. By restricting the ensemble average to these windows, CSS mitigates effects of non-uniform source distribution and restores causal–acausal symmetry in the retrieved interstation response. Applications to regional ambient-noise datasets show that CSS stabilizes surface-wave dispersion measurements even when source statistics violate assumptions of standard seismic interferometry. For the central California dataset, CSS-derived group-velocity tomograms consistently image a high-velocity block between the Rinconada and San Andreas faults across multiple periods. In comparison, the full-ensemble (linear) average does not capture this block, which is well established. Our approach is particularly useful for short-duration passive surveys.
@@ 1 Introduction
Surface-wave tomography has long been a cornerstone of the imaging of the Earth’s crust and upper mantle. The resolution of traditional surface-wave tomography, which is based on seismic waves generated by earthquakes recorded by regional and global seismic networks, is fundamentally constrained by the spatial and temporal distribution of seismicity (Trampert and Woodhouse, 2003; Ritzwoller and Levshin, 1998; Ritzwoller et al., 2001; Shapiro and Ritzwoller, 2002; Ekström, 2011; Mordret et al., 2013). Over the last 20 years, ambient seismic noise cross-correlations have supplemented surface wave tomography (Shapiro and Campillo, 2004; Sabra et al., 2005; Yang et al., 2007; Lin et al., 2008; Yang et al., 2008; Yao and Van Der Hilst, 2009; Ritzwoller et al., 2011).
@@
of $M_{ij}$ time windows provides an estimator of expectation $\mathbb{E}[\mathbf{w}_{ij}]$. Many authors have demonstrated that, in theory, the empirical average in Equation 2 converges to the interstation Green’s function, provided the ambient wavefield is isotropic, uncorrelated, and equipartitioned (Campillo and Paul, 2003; Wapenaar, 2004; Wapenaar and Fokkema, 2006; Snieder and Wapenaar, 2010). Another important theoretical result is that, when the above conditions are satisfied, $\widehat{\mathbb{E}}\left[{\mathbf{w}}_{ij}\right]$ should exhibit causal–acausal symmetry (Wapenaar, 2004; Snieder, 2004), representing equivalence between waves traveling in both directions between the receivers. In practice, however, this symmetry is rarely achieved (Bensen et al., 2007; Stehly et al., 2006; Yao and Van Der Hilst, 2009; Tsai, 2009) because real ambient noise sources are not isotropic (as shown in Figure 1a) with seasonal variations in their distribution and bursts of body wave energy (Stehly et al., 2007; Zhan et al., 2013; Delaney et al., 2017).
@@
where $\mathrm{concat}(\cdot)$ denotes the concatenation of two or more vectors, and $\mathrm{rev}(\cdot)$ denotes the reversal of the order of vector elements. Here, $\prescript{+}{}{\mathbf{w}}_{ij}$ and $\prescript{-}{}{\mathbf{w}}_{ij}$ denote the vectors associated with the causal and acausal branches of the cross-correlation, respectively, each arranged in order of increasing lag magnitude. The lack of symmetry between the causal and acausal branches means that $\widehat{\mathbb{E}}\left[\prescript{+}{}{\mathbf{w}}_{ij}\right]\neq\widehat{\mathbb{E}}\left[\prescript{-}{}{\mathbf{w}}_{ij}\right]$. If this is the case, then interstation tomography using either of the branches is not justified, since the assumption of equipartitioning is violated. In such cases, the empirical averages are also sensitive to the source distribution and more complicated sensitivity kernels have to the used during tomography (Fichtner, 2015; Ermert et al., 2017), which is computationally expensive.
@@
In this paper, we present a data-driven approach, termed Coherent Source Subsampling (CSS), which subsamples the set $\left\{\mathbf{w}_{ij}^{(m)}\right\}$ based on the state of ambient noise sources. We denote the source state using a discrete variable $\lambda$, and two such source states are illustrated in Figure 1a, using dotted and dashed lines. Our workflow for a given branch $b\in\{+,-\}$, receiver pair $(i,j)$ and source state $\lambda$ is very simple:
@@
Here, $\left|\cdot\right|$ denotes the cardinality of a set, and note that the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is not only pair specific but also branch specific, meaning the sources associated with a particular source state differ between causal and acausal branches (see Figure 1a). A straightforward approach to conditional averaging is to partition the data according to observable external variables, such as season, month, or time of day. For example, separate averages for winter and summer months could be computed, expecting coherent cross-correlations due to ocean activity or storms within a season (Yang and Ritzwoller, 2008; Zhan et al., 2013). However, such conditioning strategies may not necessarily align with the actual coherence states in the ambient wavefield.
We adopt the representation learning framework to identify coherent source states in a data-driven manner (as shown in Figure 1b). Specifically, we use a variational autoencoder to identify the subset of windows associated with a particular source state $\lambda$, rather than on predefined external variables. This idea is demonstrated using synthetic experiments for non-uniform source distributions with varying noise levels in the Supplementary Text S1. As an example of the source state, consider the stationary-phase zone for a given branch, where the sources coherently contribute to the inter-station response. In this paper, we label the source state for this zone as $\lambda=1$. Another source state is illustrated in Figure 1b. We show that CSS effectively restores causal–acausal symmetry when averaging cross-correlations confined to a particular source state. Specifically, we achieve
@@
for a choice of source state $\lambda$ and station pair $(i,j)$. Although the causal and acausal branches are redundant in the sense that they contain the same information about the interstation Green’s function, the achievement of causal–acausal symmetry is important for two reasons. First, it validates the Green’s function approximation, which is a prerequisite for robust dispersion and travel-time measurements for tomography. Second, symmetry ensures that there is no contamination from directional source effects and other transients. To our knowledge, this is the first work that achieves the theoretically predicted high degree of causal–acausal symmetry in ambient noise cross-correlations even when the underlying source distribution is highly non-uniform and time-varying. We further show that the improved symmetry leads to more accurate surface-wave tomographic maps. Finally, note that while CSS performs nonlinear averaging, similar to phase-weighted stacking by  Schimmel and Paulssen (1997), it is unique in the sense that it achieves causal–acausal symmetry.
@@ 2.1 Noise crosscorrelations
We observed that the choice of the study region is less important to the CSS framework, which applies to any ambient noise dataset. Here, we chose to demonstrate the method by analyzing continuous ambient seismic data recorded in central California using $20$ stations from the Central California Seismic Experiment (CCSE) Transportable Array deployment (CCSE, 2015). This area poses a stringent challenge for the method because the ambient-noise wavefield is highly heterogeneous and strongly directional. The Pacific Ocean acts as a dominant and persistent source of microseismic energy, leading to a marked causal–acausal asymmetry in interstation correlations (Retailleau and Beroza, 2021). In conventional ambient-noise tomography, the correlation branch associated with waves arriving from the Pacific is often preferentially used for dispersion measurements because the opposite branch is weakly illuminated (Jiang et al., 2018). The selected stations form a west–east transect that extends from the Southern Coast Ranges (SCR), through the Rinconada (RF) and San Andreas (SAF) faults system, and into the Great Valley Basin (GVB) (Supplementary Figure S55). Data were processed using standard ambient-noise interferometry procedures discussed in Bensen et al. (2007). The detailed preprocessing steps are described in the Supplementary Text S22.
Before detailing the network architecture and subsampling procedure, we first provide a simple explanation of stationary zones, highlighting the limitation of the linear empirical average in Equation 2 and motivating the conditional average in Equation 4.
@@ 2.2 Stationary-zone source state
Although the cross-correlations are coherent within stationary zones, in practice, the energy contribution from such zones does not necessarily dominate the ensemble average over all available windows (Equation 2) — this is because the source density $\rho(\mathbf{s},\mathbf{x}_{s})$ may be negligibly small in those regions. In other words, as Equation 7 indicates that the resulting distribution $\rho(\tau_{ij},\mathbf{s})$ depends jointly on the mapping $T_{ij}(\mathbf{x}_{s})$, and on the spatial source distribution $\rho(\mathbf{x}_{s},\mathbf{s})$, the departure of $\rho(\mathbf{x}_{s},\mathbf{s})$ from theoretical assumptions biases the interstation Green’s function retrieval. In such cases, the retrieval of Green’s functions can be improved by isolating and averaging only windows from stationary zones. We fix $\lambda=1$ to denote the source state of the stationary zone, which is the only state of interest in this study. Additional source states may exist, namely alternative configurations of the ambient noise field, defined by sources with similar signatures that yield coherent cross-correlations between receiver pairs for a given arrival; we label these by $\lambda=2,\ldots,K$. Figure 1a and the synthetic experiments (Supplementary Text S11) provide representative examples of such states.
@@ 2.3 Latent variable modeling
To enable data-driven identification of coherent source subsets $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$, we adopt a latent-variable modeling approach based on variational autoencoders (VAEs). VAEs provide a probabilistic framework for learning interpretable latent representations (Kingma and Welling, 2013; Doersch, 2016; Bishop and Bishop, 2023; Prince, 2023) from high-dimensional data. Within this framework, we introduce a discrete latent variable $\lambda$ (Morton et al., 2021) to represent source states and design the network architecture such that $\lambda$ correlates with physically meaningful source states, such as the stationary-zone state. The design is based on deliberate architectural decisions: a permutation-invariant encoder that accumulates coherent information (Bharadwaj, 2024) across all windows for a given pair and a categorical posterior distribution over $\lambda$ that naturally groups coherent cross-correlation windows. We emphasize that alternative network designs may exist; however, the focus of this paper is not to exhaustively explore all possible architectures, but rather to demonstrate that a carefully designed VAE can successfully solve the CSS problem by learning to distinguish stationary-zone windows from incoherent noise. The following subsections brief the generative and inference models that implement this approach.
@@ 2.3.1 Generative model
We begin with a generative model that mathematically describes how observed cross-correlations are synthesized from latent variables. That is to say, we train a decoder network that generates each cross-correlation waveform $\prescript{b}{}{\mathbf{w}}_{ij}^{(m)}$ using three distinct latent-variable components — they are highlighted in Figure 1b.
@@
Here, $l$ denotes the source state with the highest posterior probability for the window $m$. The $\mathtt{decoder}$ effectively selects a single coherent feature $\prescript{b}{l}{\mathbf{g}}_{ij}$ for generation of the observed cross-correlation. We implement it as a convolutional neural network.
@@ 2.3.2 Inference model
Here, $\mathtt{enc}_{\mathbf{g}}$, $\mathtt{enc}_{\pi}$, and $\mathtt{enc}_{\mathbf{n}}$ are convolutional encoder components that infer coherent source features, source state probabilities, and nuisance features, respectively. The symmetric encoder component $\mathtt{enc}_{\mathbf{g}}$ is permutation invariant, i.e., its output $\prescript{b}{\lambda}{\mathbf{g}}_{ij}$ is invariant to the ordering of input windows. This component accumulates information across all time windows for a given station pair and branch. On the contrary, $\mathtt{enc}_{\pi}$ and $\mathtt{enc}_{\mathbf{n}}$ operate on individual windows to capture window-specific features, inferring $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ and $\prescript{b}{}{\mathbf{n}}^{(m)}_{ij}$, respectively. Note that the autoencoder may assign the stationary-zone state to any index in $\{1,\cdots,K\}$. After training, we therefore relabeled the source states so that the stationary-zone state is always set to $\lambda=1$. Additional architectural and training details are given in the Supplementary Text S33.
@@ 2.4 Coherent source subsampling (CSS)
where the condition isolates contributions from sources associated with the state indexed by $\lambda$. It uses the learned source state probabilities $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$ for subsampling. Specifically, for each branch $b$, station pair $(i,j)$, and source state $\lambda$, the CSS estimator of the conditional expectation in Equation 11 is given by Equation 4, where the subset of correlation windows $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ is defined as:
@@
It is important to note that the identification of windows belonging to a specific source state is accomplished by probability thresholding (we used $\alpha=0.85$). By averaging only the time windows with high posterior probability $\prescript{b}{\lambda}{\pi}_{ij}^{(m)}$, we in effect subsample the windows associated with a given source state. With this targeted subsampling and by setting $\lambda=1$, we can isolate the contributions from stationary regions in the source field. Note that the conditional average $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda\right]$ in Equation 4 is a non‑linear function of the data, unlike linear averaging, because the subset $\prescript{b}{\lambda}{\mathcal{C}}_{ij}$ identification itself depends on the data. It is important to emphasize that the averaging is performed on the original cross-correlation windows, using the VAE only to identify which windows belong to each source state. Apart from the source state probabilities, we do not use any other outputs of the encoder or decoder networks, such as the reconstructed cross-correlations, in our analysis. We now demonstrate the advantages of subsampling over indiscriminate averaging.